# Harmony in 10-TET: From Parallel Universes to the Desargues Configuration

### v1

@@ 9.6. Identification with the Generalized Petersen Graph
The Levi graph of the (5,2) system is isomorphic to the Decagonal Prism Graph, which corresponds to the Generalized Petersen Graph $GP(10,1)$.
@@
The Generalized Petersen Graph $GP(n,k)$ consists of vertices $\{u_{0},\dots,u_{n-1}\}\cup\{v_{0},\dots,v_{n-1}\}$ with edges:
@@
    $v_{i}\sim v_{i+k}$ (Inner rim).
In our system, we identified two disjoint 10-cycles generated by the $P$ and $R$ transformations: the Even Universe and the Odd Universe. Let the vertices of the Even Universe be mapped to the outer rim $u_{i}$, and the vertices of the Odd Universe be mapped to the inner rim $v_{i}$, such that the $L$-transformation corresponds to the spokes connecting them.
Specifically, the $L$-transformation connects $D_{r}$ (Even) to $M_{r+5}$ (Odd). If we align the indices such that the $L$-connection is vertical ($u_{i}\sim v_{i}$), the connectivity within the Odd universe (inner rim) mirrors the connectivity of the Even universe (outer rim) with a step size of $k=1$. Thus, the graph is a prism graph $Y_{10}$, which is denoted as $GP(10,1)$. ∎
10. Isomorphism of the Acoustic and Tritone Systems
We have arrived at a remarkable theoretical unification. In Section 7, we established that the “Restored” Acoustic system (4,3)—formed by connecting two disjoint chromatic universes via “wormhole” $R$-transformations—possesses the topology of the Decagonal Prism $GP(10,1)$. Similarly, in Section 9, we proved that the Tritone system (5,2)—naturally connected by $L$-transformations—also forms the Decagonal Prism $GP(10,1)$.
This implies that despite their radically different intervallic generators ($t=4,s=3$ versus $t=5,s=2$) and different consonance criteria, these two harmonic systems are isomorphic. They describe the exact same pathway through the space of triads, merely labelled with different pitch-classes.
10.1. The Isomorphism Theorem
To formalize this equivalence, we construct a mapping $\Phi$ that transforms the chord space of the Tritone system into the chord space of the Acoustic system. The key insight is that the topological role of the transformations is swapped:
  • The Rim of the prism in (5,2) is generated by $P$ and $R$. In (4,3), the Rim is generated by $P$ and $L$.
  • The Spokes of the prism in (5,2) are the $L$ (tritone) steps. In (4,3), the Spokes are the $R$ (wormhole) steps.
Theorem 8 (Isomorphism of 10-TET Systems).
Let $\mathcal{G}_{(5,2)}$ be the Levi graph of the Tritone system and $\mathcal{G}_{(4,3)}$ be the Levi graph of the Restored Acoustic system. There exists a graph isomorphism $\Phi:\mathcal{G}_{(5,2)}\to\mathcal{G}_{(4,3)}$ defined by the index mapping $\sigma(k)=7k\pmod{10}$ acting on the chord roots, such that:
(14) $\Phi(D_{k}^{(5,2)})$ $=D_{\sigma(k)}^{(4,3)}$
(15) $\Phi(M_{k}^{(5,2)})$ $=M_{\sigma(k)}^{(4,3)}$
This mapping induces a transformation of the harmonic operations:
(16) $\Phi(P_{(5,2)})=P_{(4,3)},\quad\Phi(R_{(5,2)})=L_{(4,3)},\quad\Phi(L_{(5,2)})=R_{(4,3)}.$
@@ 10.1. The Isomorphism Theorem
The cycle of roots in the Even Universe of the (5,2) system proceeds by steps of +2+2 (via the $R$ operation). The cycle of roots in the First Universe of the (4,3) system proceeds by steps of +4+4 (via the $L$ operation). The mapping $\sigma(k)=7k$ maps the step generator $2$ to $4$, since:
$\sigma(k+2)=7(k+2)=7k+14\equiv 7k+4\pmod{10}$
Thus, adjacency along the rims is preserved. Similarly, the “Spoke” operation $L$ in (5,2) connects $D_{k}$ to $M_{k+5}$. The mapping $\sigma$ transforms this to a connection between $D_{7k}$ and $M_{7(k+5)}=M_{7k+35}\equiv M_{7k+5}$. In the (4,3) system, the $R$ operation connects chords with interval difference 5 (antipodal). Thus, the spokes are also mapped to spokes. Since $\Phi$ is a bijection preserving all edges, the graphs are isomorphic. ∎
10.2. The Double Grand Tour
The existence of this isomorphism allows us to translate any musical progression from one system to the other. In Figure 19, we illustrate the Hamiltonian Cycle found in Section 9 (The Tritone Tour) and its direct image under $\Phi$ in the Acoustic System.
For visual clarity, both graphs are presented on the standard circular Levi arrangement ($D_{0},M_{0},D_{1},M_{1}\dots$). The path on the left is the “Tritone Tour”. The path on the right is the isomorphic “Acoustic Tour”. The correspondence is indicated by the step numbers: the transition labeled $k$ on the left maps to the transition $\Phi(k)$ on the right.
$D_{0}$$M_{0}$$D_{1}$$M_{1}$$D_{2}$$M_{2}$$D_{3}$$M_{3}$$D_{4}$$M_{4}$$D_{5}$$M_{5}$$D_{6}$$M_{6}$$D_{7}$$M_{7}$$D_{8}$$M_{8}$$D_{9}$$M_{9}$1234567891011121314151617181920(a) System (5,2)
$D_{0}$$M_{0}$$D_{1}$$M_{1}$$D_{2}$$M_{2}$$D_{3}$$M_{3}$$D_{4}$$M_{4}$$D_{5}$$M_{5}$$D_{6}$$M_{6}$$D_{7}$$M_{7}$$D_{8}$$M_{8}$$D_{9}$$M_{9}$$\Phi_{1}$$\Phi_{2}$$\Phi_{3}$$\Phi_{4}$$\Phi_{5}$$\Phi_{6}$$\Phi_{7}$$\Phi_{8}$$\Phi_{9}$$\Phi_{10}$$\Phi_{11}$$\Phi_{12}$$\Phi_{13}$$\Phi_{14}$$\Phi_{15}$$\Phi_{16}$$\Phi_{17}$$\Phi_{18}$$\Phi_{19}$$\Phi_{20}$(b) System (4,3)
Figure 19. Visualizing the Parallel Grand Tours. Both systems are represented on the standard chromatic circle. Left: The Hamiltonian cycle in the (5,2) system. Right: The isomorphic cycle in the (4,3) system. The isomorphism $\Phi$ maps the Tritone Bridges (Orange, left) to the Wormhole Connections (Violet, right).
10.3. Musical Interpretation
The formal isomorphism reveals a profound duality between the two systems. In the Tritone System (5,2), the “difficult” interval is the Tritone ($t=5$). It is the interval that forces a jump between the Even and Odd parities. In the Acoustic System (4,3), the “difficult” interval is the Wormhole (antipodal connection). It is the interval that forces a jump between the two disjoint hexatonic-like universes.
Mathematically, Parity $\cong$ Universe. Navigating between Even and Odd chords using tritones in (5,2) is structurally indistinguishable from navigating between parallel harmonic universes using wormholes in (4,3). The “Grand Tour” of the 10-TET scale is thus a universal invariant: whether one views it through the lens of (5,2) or (4,3), the path requires exactly one bridge-crossing (and one return) to visit every harmony.
@@ Acknowledgements
Appendix A Appendix A: Historical Motivation and the Acoustic Implementation

### v2

@@ 10.1.1. The “Acoustic” system $(4,3)$
The functional 10-TET Acoustic graph is isomorphic to the Bipartite Circulant Graph $Ci_{20}(1,9,15)$.
@@
To analyze the global structure, we map the 20 distinct chords into the cyclic group $\mathbb{Z}_{20}$. We arrange the bipartite sets in an alternating (interleaved) fashion using the mapping $\phi:\mathcal{C}\to\mathbb{Z}_{20}$:
$\phi(D_{i})=2i\quad\text{and}\quad\phi(M_{i})=2i+1\pmod{20}$.
Here $\mathcal{C}=\{D_{i},M_{j},\,\,i,j=0,1,\dots 9\}$ is our musical space of all chords. In this unified space, an edge connects a Major chord $D_{i}$ to a Minor chord $M_{i+k}$ (where $k$ is the modular offset determined in Section 6.1). The geometric “jump” or distance $J$ between these vertices in $\mathbb{Z}_{20}$ is:
$J$ $=\phi(M_{i+k})-\phi(D_{i})$
$=(2(i+k)+1)-2i$
$=2k+1\pmod{20}$
Using the receipt for the Major/Minor chords assocciation $D_{r}\to(M_{r},M_{r+4},M_{r+7})$ for the Acoustic system as in Section 7.2.1, we get the onnection offsets as $k\in\{0,4,7\}$, and calculate the raw jumps:
@@
    $k=7\implies J=15$.
These are three generators $s$ of the corresponding Circulant graph. Since the graph is vertex-transitive it is the circulant graph $Ci_{20}(1,9,15)$. ∎
10.1.2. The “Tritone” system
Theorem 7.
The functional 10-TET Tritone graph is isomorphic to the Bipartite Circulant Graph $Ci_{20}(1,11,17)$.
@@ 10.1.2. The “Tritone” system
We employ the same bipartite mapping $\phi:\mathcal{C}\to\mathbb{Z}_{20}$ used in the proof for the Acoustic system:
$\phi(D_{i})=2i\quad\text{and}\quad\phi(M_{i})=2i+1\pmod{20}$
This mapping allows us to determine the graph structure by calculating the geometric distance (or “jump”) between connected vertices. For a connection between $D_{i}$ and $M_{i+k}$, the distance in $\mathbb{Z}_{20}$ is given by $J=2k+1$.
The Tritone system is defined by the connections where each Major chord $D_{i}$ meets the Minor chords $M_{i}$, $M_{i+5}$, and $M_{i+8}$. This establishes the offset set $S=\{0,5,8\}$. We calculate the generators for the circulant graph as follows:
  1. (1)
    Offset $k=0$: The distance is $2(0)+1=\mathbf{1}$.
  2. (2)
    Offset $k=5$: The distance is $2(5)+1=\mathbf{11}$.
  3. (3)
    Offset $k=8$: The distance is $2(8)+1=\mathbf{17}$.
The resulting set of symmetric generators is $\{1,11,17\}$. Thus, the graph is exactly the bipartite circulant graph $Ci_{20}(1,11,17)$. ∎
10.1.3. The “Wide” system $(6,1)$
We end this analysis by also viewing the graph of the “Wide” system $10_{3}$ configuration as a circulant. This, as we see below, will independently confirm our earlier identification of this system as the cyclic $(10_{3})_{3}$ configuration of [3].
Theorem 8.
The functional 10-TET Wide graph is isomorphic to the Bipartite Circulant Graph $Ci_{20}(1,13,19)$, which constitutes the Hamiltonian representation of the unique cyclic configuration $(10_{3})_{3}$.
Proof.
We apply the standard bipartite mapping $\phi:\mathcal{C}\to\mathbb{Z}_{20}$ to the Wide system. The connection offsets for this system, as derived in Section 8.6, are $S=\{0,6,9\}$. We calculate the corresponding geometric generators $J=2k+1$ in the cyclic graph:
  1. (1)
    Offset $k=0$: The jump is $2(0)+1=\mathbf{1}$.
  2. (2)
    Offset $k=6$: The jump is $2(6)+1=\mathbf{13}$.
  3. (3)
    Offset $k=9$: The jump is $2(9)+1=\mathbf{19}$.
The resulting set of generators is $\{1,13,19\}$. Note that the generators $1$ and $19$ (which is 1-1) correspond to the two directions of the same edges along the perimeter of the cycle. This indicates that the graph consists of a Hamiltonian cycle $C_{20}$ (formed by the $\pm 1$ edges) supplemented by chords of span 13 connecting every vertex $v$ to $v+13$ (or equivalently v7v-7 in the bipartite mapping).
Since this specific circulant describes a $10_{3}$ configuration, and since by its very contruction it is cyclic, it must be identified with the only cyclic $10_{3}$ configuration present in census [3]. Thus, we confirm that the Wide system is the system of $(10_{3})_{3}$ configuration from the classification list in [3], p.335. ∎
10.2. Automorphism groups
We end this article by coments on symmetry groups of the 10-TET harmony systems. They are based on our calculations using computational graph theory Python libraries.
We start by defining a musical system automorphism group, as an authomorphism group of its corresponding graph. The latter is defined as follows:
Definition 2 (Graph Automorphism and Geometric Symmetry).
Let $G=(V,E)$ be a graph with set of points $V$ and set of edges $E$. An automorphism of $G$ is a permutation $\sigma:V\to V$ such that for any pair of vertices $u,v\in V$, $\{u,v\}\in E$ if and only if $\{\sigma(u),\sigma(v)\}\in E$. The set of all such automorphisms forms the Geometric Symmetry Group of $G$, denoted $\text{Aut}(G)$.
Within this definition, our calculations reveal that:
  • The “Acoustic” $(4,3)$ solution system has group of authomorphisms of order 40.
  • The “Tritone” $(5,2)$ solution system has group of authomorphisms of order 320. It is large, roughly because of the existence of the “Tritone” jump $k\to k+5$, with $5=\frac{n}{2}$.
  • The “Wide” $(6,1)$ solution system has group of authomorphisms of order 20, isomorphic – via vivible symmetry of its circulant graph – to the dihedral group $D_{10}$. This agrees with the symmetry group of the cyclic $(10_{3})_{3}$ confifguration, and confirms our earlier identification of this system at the end of Theorem 8.
@@ Acknowledgements
Appendix A Historical Motivation and the Acoustic Implementation